sqrt(2p-1) + sqrt(p-1)=1
Answer by stanbon(75887)

sqrt(2p-1) + sqrt(p-1)=1
-----
Square both sides to get:
2p-1 + p-1 + 2sqrt[(p-1)(2p-2)] = 1
----------------------------
3p-3 = 2sqrt[(p-1)(2p-2)]
---------
Square both sides:
9p^2 - 18p + 9 = 4(p-1)(2p-2)
-----
9p^2 -18p + 9 = 8(p^2-3p+2)
9p^2 -18p + 9 = 8p^2 - 24p + 16
p^2 + 6p -7 = 0
-----
Factor:
(p+7)(p-1) = 0
----
Possible solutions: p = 1 or p = -7
-------
Check each of the possible solutions in the original equation.
==================
Cheers,
Stan H.